Shift Summary for Evening shift, 05/27/99

Jianbo Wang (jianbo@fnal.gov)
Thu, 27 May 1999 23:29:20 -0500 (CDT)

Shift Summary for Evening shift, 05/27/99

On Shift: M Arenton
J Wang
People in the control room: BobT, Ed, Taku, RickK, Sydney, JasonH, JasonL,
Jim, Sandrine, RickT, Sasha, Elizabeth, JAWs, Nick, Hogen, Miguel

A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0

Accounting of lost time:

To test HCC, HCC bit threshold will be lowered to het turned on by MIPS.
To calibrate between DAC and output voltage, Jason H remeasured 1-60
boards and put them into ETOT.config

RJT added trigger 13 to cosmic trigger configuration to monitor muon
system.
HA-trigger 2 (V*MU2)
MU2,MU3- trigger 2,13 (MU2*(MU3x+MU3Y))

Dot could not append to the end of tape when mounted. the underlying ftt
calls use the raw SCSI devices (/dev/scsi) to position the tape.
Permissions on these files were such that only root could rw (644). They
need write access for all users(666). DON'T SET THIS FOR ALL SCSI DEVICES.
You don't want users to accidentally open disk drives in raw scsi mode.

Taku removed "-d" option from the 1st line, updated run control
to V3.09 so that it set up correct DIS parameters for cache file size.

Arun changed the path ~da_ktev/dart/config/ktev97/targets/ktev3.csh to
damp-summary v1-8

Bob T and Jianbo walk around the hall and check two magnets and Z-cal.
Z-cal is in-position for muon run. The block house is locked.

Etot thresholds returned to normal.

Sydney loads ETOT in run control and KTRIG.

People from main control room should come at 8:30pm to do security check.
But hey never showed up. There is no beam tonight.

========================================================================
Runs ended this shift:
Run 13146 - completed
Run 13147 - cosmic run (still running at the end of the shift)

Tapes brought to FCC this shift: 0